Suppose T "\isin" L(V) is invertible.

(a) Suppose "\lambda\isin" F with "\lambda" not= 0. Prove that "\lambda" is an eigenvalue of T if and only if 1/"\lambda" is an eigenvalue of T-1.

(b) Prove that T and T-1 have the same eigenvectors.


1
Expert's answer
2021-07-07T11:14:25-0400

(a)

"\lambda" is an eigenvalue of "T"

"\iff" there exists a nonzero "v\in V" such that "Tv=\lambda v"

"\iff" there exists a nonzero "v\in V" such that "T^{-1}Tv=T^{-1}\lambda v"

"\iff" there exists a nonzero "v\in V" such that "v=\lambda T^{-1}v"

"\iff" there exists a nonzero "v\in V" such that "\dfrac{1}{\lambda}v=T^{-1}v"

"\iff \dfrac{1}{\lambda}" is an eigenvalue of "T^{-1}."


(b)

Note that since "T" is invertible, "T" is injective.

Suppose "(0, v)" is an eigencouple of "T." Then "Tv=0" ; thus "v=0" and in fact there cannot be any eigenvectors of "T" corresponding to "0."

Now, suppose "v" is an eigenvector of "T" corresponding to "\lambda\not=0."

By the proof of (a), "v" is an eigenvector of "T^{-1}" corresponding to "\dfrac{1}{\lambda}."

So all eigenvectors of "T" are eigenvectors of "T^{-1}"; reversing the roles of "T" and "T^{-1}"yields the reverse inclusion and completes the proof.